Kathy and Ken work in a textbook warehouse. Kathy can fill an order of 100 books in 40 minutes. Ken can fill the same order in 60 minutes. How long would it take them, working together, to fill the order?
Answer by vleith(2983)

Kathy can do 100/40 books per minute = 2.5
Ken can do 100/60 books per minutes = 1.666
together they can do 4.1666 books per minute
To do 100 books, they would need 100/4.1666 = 24 minutes